The body with mass 20kg moves in a straight line uniformly accelerated with acceleration 2m / s², on a horizontal surface, under the action of force 50N Find: a) What is the force of friction acting on the body. b) The coefficient of friction between the body and the surface where it moves.


1
Expert's answer
2021-02-08T18:42:39-0500

a) Applying the Newton's Second Law of Motion, we get:


"F-F_{fr}=ma,""F_{fr}=F-ma=50\ N-20\ kg\cdot2\ \dfrac{m}{s^2}=10\ N."

b)

"F_{fr}=\mu mg,""\mu=\dfrac{F_{fr}}{mg}=\dfrac{10\ N}{20\ kg\cdot9.8\ \dfrac{m}{s^2}}=0.05."